together a computer, monitor, and printer cost 1775. The monitor cost 300 less than the computer, and the printer cost 50 more than the monitor. Find the cost of the printer.
Found 2 solutions by mananth, nyc_function:
Answer by mananth(16946)

together a computer, monitor, and printer cost 1775. The monitor cost 300 less than the computer, and the printer cost 50 more than the monitor. Find the cost of the printer.
..
let cost of computer be $x
monitor cost will be x-300
printer cost will be x-300+50
..
sum of all = 1775
...
x+x-300+x-300+50=1775
3x-550=1775
+550
3x=1775+550
3x=2325
/3
x= 775
printer will cost 775-300+50
$ 525
m.ananth@hotmail.ca



Answer by nyc_function(2741)

Let x = computer

Let (x - 300) = monitor

Let (x + 50) = printer

x + x - 300 + x + 50 = 1775

Can you finish?